In an ideal Vapor Compression Refrigeration Cycle,using h2=382 kJ/kg, h3=243.15 kJ/kg, h4=243.15 kJ/kg and h1=338.14 kJ/kg, determine the refrigerating effect in kJ/kg and coefficient of performance.
Answer #1

h1 = 338.14 KJ/Kg

h2 = 382 KJ/Kg

h3 = h4 = 243.15 KJ/Kg

Refrigeration effect (RE) = h1 - h4 = 338.14 - 243.15 = 94.99 KJ/Kg

Work Input to the system (W) = h2 - h1= 382 - 338.14 = 43.86 KJ/Kg

COP = RE / W = 94.99 / 43.86 = 2.166